OpenStudy (anonymous):

pythagorean identity to find sin t cos t = -1/2 and pi

OpenStudy (amistre64):

s^2 + c^2 = 1 ... solve for s

OpenStudy (amistre64):

keep in mind the quadrant its asking for, as that tells us whether or not sin is positive or negative

OpenStudy (anonymous):

well I did sint^2=1+(1/2)^2 sin t = 1+1/4 sin t = -sqrt 5/2 but it's the wrong answer =/

OpenStudy (amistre64):

of course its wrong .. s^2 + c^2 = 1 - c^2 - c^2 ---------------- s^2 = 1-c^2 ; not 1+c^2

OpenStudy (anonymous):

cost is a negative, negative and negative is a positive..

OpenStudy (amistre64):

doesnt matter ....

OpenStudy (anonymous):

well if subtracting cost wouldn't it be 1- (-1/2)^2 ?

OpenStudy (amistre64):

s^2 + (-1/2)^2 = 1 s^2 + 1/4 = 1 - 1/4 - 1/4 ----------------- s^2 = 1 - 1/4

OpenStudy (anonymous):

oh um.... hold on let see....

OpenStudy (anonymous):

Oh I see, thanks! I see my mistake, can you help me with just 1 more same concept got the wrong answer but idk why.

OpenStudy (amistre64):

sure

OpenStudy (anonymous):

cost = 1/2 and 0<t<pi/2 I got sqrt 1/2 but um its wrong =/

OpenStudy (amistre64):

quad 1, so cos and sin are + .. cos is still 1/2 .... so sin is still gonna be 1-(1/2)^2 is still gonna be sqrt(3)/2 but in quad 1 all is positive

OpenStudy (amistre64):

in other words, you didnt square the c in this case

OpenStudy (anonymous):

Ohh wow i forgot, thanks for clarifying!! I see my mistake now!

OpenStudy (amistre64):

the hardest part in math is not learning new concepts, its holding on to the stuff we already learned lol

OpenStudy (anonymous):

Lol I know i easily get confused at wrong answers lol.

OpenStudy (amistre64):

good lcuk ;)

OpenStudy (anonymous):

are you a math major in colelge ?

OpenStudy (amistre64):

luck even ....

OpenStudy (amistre64):

just graduating with a BA in math ...

OpenStudy (anonymous):

Oh wow what college ?

OpenStudy (amistre64):

saint leo university ... it was the closest place to where i live.

OpenStudy (anonymous):

Oh nice, well thanks again for your help, im so stressing for my finals next week lol got a lot to cover.

OpenStudy (amistre64):

youre welcome, and relax, youll do fine

OpenStudy (anonymous):

aw thanks lol.